Describe a situation that may cause manufacturing overhead to be over-allocated in a given year. Also, describe a situation that may cause manufacturing overhead to be under-allocated in a given year.
Answer #1
Describe a situation that may cause manufacturing overhead to be over-allocated in a given year.
ABC Company allocates overhead on the basis of machine
hours. The budgeted overhead for the last period was
$100000 and the estimated machine hours was 20000 at the
beginning of the period.
The actual machine hours used turned out to be 21000 and
the actual manufacturing overhead was $90000.
The calculations would be as under:
Pre-determined overhead rate = 100000/20000 = $5 per MH.
Actual overhead allocated = 21000*$5 = $105000
As overhead allocated is more than the actual overhead of
$90000, overhead is over-allocated to the extent of
$15000 ($105000-$90000).
Also, describe a situation that may cause manufacturing overhead to be under-allocated in a given year.
For the above budget situation, if, the actuals are say;
Machine hours used were 20000 and
the actual manufacturing overhead was $110000.
The calculations would be as under:
Pre-determined overhead rate = 100000/20000 = $5 per MH.
Actual overhead allocated = 20000*$5 = $100000
As overhead allocated is less than the actual overhead of
$110000, overhead is under-allocated to the extent of
$10000 ($110000-$100000).